# Orientation Matters: Making 3D Generative Models Orientation-Aligned

### v1

@@ 5.1 Zero-shot Model-free Object Orientation Estimation
One popular class of solutions to 3D object orientation estimation is based on analysis-by-synthesis [46, 3, 10], but most of them are limited to category-level due to the difficulty in synthesizing orientation-aligned 3D models across categories. Another line of methods [28, 53] estimates poses relative to a given CAD model, limiting practical use since a corresponding model must be available for each input image. While one could replace CAD models with generated 3D shapes (e.g., from Trellis), the misaligned outputs of existing generative models result in unreliable pose references. In contrast, our method generates orientation-aligned 3D shapes directly from a single image, enabling the prediction of absolute poses in a canonical frame by treating the generated 3D shape as a per-object template. Specifically, we build on a state-of-the-art template-based pose estimation method, FoundationPose [53], which has strong generalizability due to training on a large synthetic dataset. It renders templates from a fixed set of viewpoints, refines each pose, and then selects the best match. Although FoundationPose is trained with accurate CAD models and depth maps, we find that its pose refinement module remains effective even with our generated 3D shapes and without depth input. However, its pose selection performance degrades when templates differ slightly in geometry or appearance. To address this, we retain the pose refinement module but replace the pose selection stage with a DINOv2 [26]-feature-based similarity metric, as shown in Figure 4. This is achieved by computing the L2 distance between DINOv2 features of each refined rendering and the target image, and selecting the view with the highest similarity. Note that when estimating orientations of objects in scene-level images, we need to extract objects’ masks via the image segmentation method [13] before 3D generation.
@@ 5.2 Efficient Arrow-based Object Rotation Manipulation
Efficient manipulation of 3D model rotation within simulation systems is crucial, yet challenging, especially when models are initialized in non-canonical poses. This difficulty arises because users typically aim to orient a model toward a desired direction, but conventional 3D simulation systems only record the model’s pose relative to its initial state. As a result, users must manually compensate for any misalignment in the initial pose, complicating the interaction process. In contrast, our generated orientation-aligned 3D models enable a more direct and user-friendly rotation manipulation approach due to their aligned initial orientations. With this alignment, users can simply draw an arrow indicating the desired forward-facing direction, without needing to consider the model’s original pose. This arrow-based interaction paradigm enhances usability and is compatible with both augmented reality (AR) applications and general-purpose 3D software. In AR applications, users draw an arrow in the 2D image plane, which is subsequently lifted into 3D space using monocular depth estimation techniques (e.g., [30]). The system then rotates the model so that its forward-facing axis aligns with the specified arrow, while ensuring the object remains grounded on the background plane. Note that because our models are normalized in scale, their size must still be specified, either through a large language model (LLM) or direct user input. In general 3D software, users draw the arrow directly in 3D space. The system then applies Rodrigues’ rotation formula to align the model’s orientation with the user-specified direction. Additional implementation details are provided in the supplementary materials.
@@ 6 Experiment
| GSO [8] | Toys4k [38] | |||||
| CD | LPIPS | CLIP | CD | LPIPS | CLIP | |
| Wonder3D | 0.0894 | 0.2741 | 73.24 | 0.0932 | 0.2773 | 83.14 |
| Wonder3D + PCA | 0.0852 | 0.2649 | 72.27 | 0.0929 | 0.2617 | 82.01 |
| Wonder3D + VLM (Gemini-2.0) [31] | 0.0850 | 0.2703 | 73.18 | 0.0880 | 0.2717 | 83.40 |
| Wonder3D + Orient Anything (ViT-L) [51] | 0.1015 | 0.2665 | 72.46 | 0.1079 | 0.2660 | 81.91 |
| Wonder3D-OA (ours small) | 0.0605 | 0.2379 | 77.54 | 0.0581 | 0.2285 | 89.00 |
| Wonder3D-OA (ours) | 0.0564 | 0.2338 | 78.26 | 0.0548 | 0.2253 | 89.05 |
@@ 6.1 Implementation Details
Dataset. Orientation-aligned 3D generative models Trellis-OA and Wonder3D-OA are trained on our Objaverse-OA dataset, which is curated from Objaverse-LVIS [6]. The base multi-view diffusion model is trained on Objaverse [6], and the base 3D-VAE-based model is trained on TRELLIS-500K [56]. To demonstrate the generalizability and accuracy of our method’s orientation alignment ability, we evaluate on two unseen datasets, GSO [8] and Toys4k [38]. To further demonstrate the sim-to-real generalizability, we also evaluate on the real-world dataset Imagenet3D [23].
Baselines. For the task of aligned object generation, there are no existing baselines for this task. Therefore, we design baselines that perform this task in two stages: 1) object generation with misaligned orientations, and 2) orient them to aligned poses based on pose estimation using different variants: (i) Principal Component Analysis (PCA); (ii) advanced Vision Language Model (VLM) Gemini-2.0 [31]; and (iii) zero-shot model-free orientation estimation method, Orient Anything [51]. For the task of zero-shot orientation estimation, we compare our method with Orient Anything [51] and FSDetView [57]. Note that FSDetView doesn’t support zero-shot estimation. Therefore, we evaluate its performance only on its supported categories.
Metrics. To evaluate the orientation alignment ability, we rotate reconstructed 3D models using different kinds of methods and calculate Chamfer Distance (CD), LPIPS [62], and CLIP [33] scores to measure the orientation alignment quality. To evaluate the performance of our zero-shot orientation estimation method, we calculate Acc@30 and orientation absolute error (Abs) according to the rotation error. We follow NOCS to calculate the rotation defined by: , where represents the trace of the matrix. Note that for stick-like objects, top and side directions typically have ambiguity. Therefore, we only calculate the rotation error in the front direction.
Training and inference time. To fine-tune Trellis-OA, we use a total batch size of 64 for training 30000 steps, which takes only about 10 hours on the cluster of 8 Nvidia Tesla A100 GPUs. To fine-tune Wonder3D-OA, we use a total batch size of 512 for training 40000 steps, which takes about 3 days on the cluster of 8 Nvidia Tesla A100 GPUs.. For the 3D object generation, Trellis-OA takes 32.93s and Wonder3D-OA takes 6.69s for multi-view generation, 2.53s for fused 3DGS generation, and 45.37s for mesh extraction. Given the generated 3D models, our pose refinement module takes 6.54s and the pose selection module takes 7.47s, which are comparable with the SOTA template-based pose estimation method FoundationPose [53].
@@ 6.2 Orientation-Aligned Object Generation
| GSO [8] | Toys4k [38] | |||||
@@
| Trellis + PCA | 0.0912 | 0.2771 | 78.84 | 0.1031 | 0.2758 | 88.09 |
| Trellis + VLM (Gemini-2.0) [31] | 0.0421 | 0.1998 | 89.97 | 0.0564 | 0.2137 | 95.19 |
| Trellis + Orient Anything (ViT-L) [51] | 0.0604 | 0.2193 | 88.61 | 0.0725 | 0.2313 | 94.81 |
@@
Comparison with baselines: We first evaluate the performance of orientation-aligned object generation. As shown in the Table 1 and Table 2, our fine-tuned models, Wonder3D-OA and Trellis-OA, surpass most two-stage baselines in geometry and appearance on both GSO and Toys4k datasets. Note that our method is suboptimal in appearance on the GSO dataset compared to the VLM baseline. It is probably because the GSO dataset has more objects with irregular geometry and appearance, which is challenging for our fine-tuned generator to handle. We present more qualitative results in the Figure 5 and Figure 6. As shown in Figure 6, PCA cannot handle objects with different shape features and has difficulty in distinguishing the direction of estimated principal axes. VLM has trouble recognizing objects with unclear front-view features, like dolphins and boats. Orient Anything is still not that accurate, and typically rotates objects into leaning poses. In contrast, our method accurately aligns the orientation of the generated 3D models without obvious geometry and appearance degradation. Besides, due to the high geometry quality of our Objaverse-OA, Trellis-OA eliminates the production of plane-like geometries. Please refer to the supplementary materials for more qualitative results.
@@ 6.3 Zero-Shot Object Orientation Estimation
| Toys4k [38] | Stick-like Obj. from ImageNet3D [23] | |||
@@
| FSDetView [57] (Few-shot) | 20.90 | 91.66 | 10.29 | 84.25 |
| Orient Anything [51] (Vit-S) | 42.05 | 52.72 | 2.63 | 81.70 |
| Orient Anything [51] (Vit-L) | 63.18 | 36.37 | 9.8 | 78.19 |
@@
Comparison with baselines: As shown in Table 3, without specific training, our method is already comparable with the SOTA orientation estimation method [51] with the ViT-large architecture and surpasses [51] with the ViT-small architecture by a large margin. Besides, our method can handle challenging stick-like objects collected from the real-world dataset while baselines all fail to work, which further demonstrates our utility in coping with long-tailed situations.
@@ 7 Conclusion and Limitation
In this appendix, we detail our dataset curation, experiment implementations, and method implementations in Section A. Besides, we provide more results of our efficient arrow-based object manipulation, orientation-aligned object generation, orientation estimation, and failure cases in Section B.
@@ A.1 Dataset Curation
Data Processing Pipeline. We first apply Vision Language Model (VLM) recognition across the entire dataset, followed by manual identification and correction of failure cases. Specifically, four images were rendered for each 3D object using an orthogonal camera setup. The front view was identified using the VLM (Gemini 2.0 [31]), and the objects were subsequently rotated to their canonical poses based on the recognition results. Note that objects recognized as without front-view orientation are automatically excluded from the dataset. After this initial step, the orientations of the processed objects were manually refined using Blender software.
@@ A.2 Experiment Implementations
Evaluation Data. We evaluate our method on three unseen datasets: GSO [8], Toys4k [38], and Imagenet3D [23]. For the GSO dataset, we randomly collected 48 objects with clearly defined front views. Each object was rendered into four images, sampled from the upper hemisphere with azimuth angles uniformly distributed within [0∘, 360∘], polar angles within [0∘, 60∘], and rotation angles within [-30∘, 30∘]. For the Toys4k dataset, we randomly selected 439 objects across 47 categories. Each object was rendered into a single image using the same camera configuration as applied to the GSO dataset. For the Imagenet3D dataset, we focused on seven stick-like object categories, comprising a total of 204 objects. Please refer to Table 4 and Table 5 for details.
@@
Baselines. To implement the PCA-based baseline, we first calculate the three principal axes by eigen-decomposing the 3D models’ geometries. The object is then rotated to align these principal axes with the global coordinate system’s x-, y-, and z-axes. For the VLM-based baseline, we follow the same method used in the dataset curation, utilizing Gemini-2.0 [31] for orientation recognition. To implement the baseline based on Orient Anything [51], we use the official checkpoint based on the ViT-large architecture and adopt its data augmentation module. The object orientation is estimated from an image rendered using a fixed camera, and the object is subsequently rotated according to the predicted orientation. For the orientation estimation baseline based on FSDetView [57], we restrict evaluation to the object categories supported by the method.
@@ A.3 Method Implementations
Efficient Arrow-based Object Rotation Manipulation. For the arrow-based object rotation manipulation in the augmented reality application, we define this arrow-based interaction using a 2D start point and a 2D end point , forming a direction vector . Then we use Unidepth [30] to estimate the camera intrinsic and depth map. With camera intrinsic, we calculate the 3D camera rays and for the corresponding 2D points and . With the depth map, we calculate the plane via the least square method. After that, we calculate the 3D intersection points and between the corresponding camera rays , and the extracted plane . Finally, 3D direction vector can be calculated: . Since each generated model is aligned to a known forward vector , we can directly compute and apply the rotation that aligns with , allowing the object to be placed in the correct orientation without any manual adjustment. Note that the 3D start point is also the 3D location of the inserted 3D object, the 3D models need to rotate along to ensure verticality to the ground and the object size still needs to be set via VLM or user interaction, since our generated models share a normalized scale. Specifically, to automatically estimate the size of the object to insert, we can directly ask the VLM to estimate the common size of the object in the real world with the single-view image as input, since our method estimates metric depth with the scale aligned with the real world for insertion. Besides, to make the insertion results realistic, we set a plane as a shadow catcher and estimate the environment lighting using DiffusionLight [29]. For the arrow-based object rotation manipulation in the generic 3D software, we define this arrow-based interaction using a 3D start point and a 3D end point , forming a 3D direction vector . The rotation transformation is calculated by aligning with using Rodrigues’ formula and users can further rotate the objects along if needed.

### v2

@@ 5.1 Zero-shot Model-free Object Orientation Estimation
One popular class of solutions to 3D object orientation estimation is based on analysis-by-synthesis [Wang2019NormalizedOC, Chen2020CategoryLO, Guo2022AVN], but most of them are limited to category-level due to the difficulty in synthesizing orientation-aligned 3D models across categories. Another line of methods [Peng2018PVNetPV, Wen2023FoundationPoseU6] estimates poses relative to a given CAD model, limiting practical use since a corresponding model must be available for each input image. While one could replace CAD models with generated 3D shapes (e.g., from Trellis), the misaligned outputs of existing generative models result in unreliable pose references. In contrast, our method generates orientation-aligned 3D shapes directly from a single image, enabling the prediction of absolute poses in a canonical frame by treating the generated 3D shape as a per-object template. Specifically, we build on a state-of-the-art template-based pose estimation method, FoundationPose [Wen2023FoundationPoseU6], which has strong generalizability due to training on a large synthetic dataset. It renders templates from a fixed set of viewpoints, refines each pose, and then selects the best match. Although FoundationPose is trained with accurate CAD models and depth maps, we find that its pose refinement module remains effective even with our generated 3D shapes and without depth input. However, its pose selection performance degrades when templates differ slightly in geometry or appearance. To address this, we retain the pose refinement module but replace the pose selection stage with a DINOv2 [Oquab2023DINOv2LR]-feature-based similarity metric, as shown in Figure 4. This is achieved by computing the L2 distance between DINOv2 patch feature maps of each refined rendering and the target image, and selecting the view with the highest similarity. Note that when estimating orientations of objects in scene-level images, we need to extract objects’ masks via the image segmentation method [Kirillov2023SegmentA] before 3D generation.
@@ 5.2 Efficient Arrow-based Object Rotation Manipulation
Efficient manipulation of 3D model rotation within simulation systems is crucial, yet challenging, especially when models are initialized in non-canonical poses. This difficulty arises because users typically aim to orient a model toward a desired direction, but conventional 3D simulation systems only record the model’s pose relative to its initial state. As a result, users must manually compensate for any misalignment in the initial pose, complicating the interaction process. In contrast, our generated orientation-aligned 3D models enable a more direct and user-friendly rotation manipulation approach due to their aligned initial orientations. With this alignment, users can simply draw an arrow indicating the desired forward-facing direction, without needing to consider the model’s original pose. This arrow-based interaction paradigm enhances usability and is compatible with both augmented reality (AR) applications and general-purpose 3D software. In AR applications, users draw an arrow in the 2D image plane, which is subsequently lifted into 3D space using monocular depth estimation techniques (e.g., [Piccinelli2024UniDepthUM]). The system then rotates the model so that its forward-facing axis aligns with the specified arrow, while ensuring the object remains grounded on the background plane. Note that because our models are normalized in scale, their size must still be specified, either through a large language model (LLM) or direct user input. In general 3D software, users draw the arrow directly in 3D space. The system then applies Rodrigues’ rotation formula to align the model’s orientation with the user-specified direction. Additional implementation details are provided in the supplementary materials Section A.3.
| GSO [Downs2022GoogleSO] | Toys4k [Stojanov2021UsingST] | |||||
| CD | LPIPS | CLIP | CD | LPIPS | CLIP | |
| Wonder3D | 0.0894 | 0.2799 | 76.37 | 0.0932 | 0.2859 | 87.10 |
| Wonder3D + PCA | 0.0788 | 0.2554 | 77.80 | 0.0858 | 0.2691 | 87.58 |
| Wonder3D + VLM (Gemini-2.0) [gemini] | 0.0850 | 0.2752 | 76.30 | 0.0880 | 0.2804 | 87.53 |
| Wonder3D + Orient Anything (ViT-L) [Wang2024OrientAL] | 0.1015 | 0.2600 | 77.50 | 0.1079 | 0.2699 | 88.12 |
| Wonder3D-OA (ours w/o LGM) | 0.0609 | 0.2300 | 80.22 | 0.0571 | 0.2351 | 91.33 |
| Wonder3D-OA (ours) | 0.0564 | 0.2270 | 80.30 | 0.0548 | 0.2317 | 92.09 |
@@ 6.1 Implementation Details
Dataset. Orientation-aligned 3D generative models Trellis-OA and Wonder3D-OA are trained on our Objaverse-OA dataset, which is curated from Objaverse-LVIS [Deitke2022ObjaverseAU]. The base multi-view diffusion model is trained on Objaverse [Deitke2022ObjaverseAU], and the base 3D-VAE-based model is trained on TRELLIS-500K [Xiang2024Structured3L]. To demonstrate the generalizability and accuracy of our method’s orientation alignment ability, we evaluate on two unseen datasets, GSO [Downs2022GoogleSO] and Toys4k [Stojanov2021UsingST]. To further demonstrate the sim-to-real generalizability, we also evaluate on the real-world dataset Imagenet3D [Ma2024ImageNet3DTG].
Baselines. For the task of aligned object generation, there are no existing baselines for this task. Therefore, we design baselines that perform this task in two stages: 1) object generation with misaligned orientations, and 2) orient them to aligned poses based on pose estimation using different variants: (i) Principal Component Analysis (PCA); (ii) advanced Vision Language Model (VLM) Gemini-2.0 [gemini]; and (iii) zero-shot model-free orientation estimation method, Orient Anything [Wang2024OrientAL]. For the task of zero-shot orientation estimation, we compare our method with Orient Anything [Wang2024OrientAL] and FSDetView [Xiao2020FewShotOD]. Note that FSDetView doesn’t support zero-shot estimation. Therefore, we evaluate its performance only on its supported categories.
Metrics. To evaluate the orientation alignment ability, we rotate reconstructed 3D models using different kinds of methods and calculate Chamfer Distance (CD), LPIPS [Zhang2018TheUE], and CLIP [CLIP2021ICML] scores to measure the orientation alignment quality. To evaluate the performance of our zero-shot orientation estimation method, we calculate Acc@30 and orientation absolute error (Abs) according to the rotation error. We follow NOCS to calculate the rotation defined by: , where represents the trace of the matrix. Note that for stick-like objects, top and side directions typically have ambiguity. Therefore, we only calculate the rotation error in the front direction.
Training and inference time. To fine-tune Trellis-OA, we use a total batch size of 64 for training 30000 steps, which takes only about 10 hours on the cluster of 8 Nvidia Tesla A100 GPUs. To fine-tune Wonder3D-OA, we use a total batch size of 512 for training 40000 steps, which takes about 3 days on the cluster of 8 Nvidia Tesla A100 GPUs.. For the 3D object generation, Trellis-OA takes 32.93s and Wonder3D-OA takes 6.69s for multi-view generation, 2.53s for fused 3DGS generation, and 45.37s for mesh extraction. Given the generated 3D models, our pose refinement module takes 6.54s and the pose selection module takes 7.47s, which are comparable with the SOTA template-based pose estimation method FoundationPose [Wen2023FoundationPoseU6].
@@ 6.2 Orientation-Aligned Object Generation
| GSO [Downs2022GoogleSO] | Toys4k [Stojanov2021UsingST] | |||||
@@
| Trellis + PCA | 0.0798 | 0.2771 | 78.84 | 0.0911 | 0.2758 | 88.09 |
| Trellis + VLM (Gemini-2.0) [gemini] | 0.0421 | 0.1998 | 89.97 | 0.0564 | 0.2137 | 95.19 |
| Trellis + Orient Anything (ViT-L) [Wang2024OrientAL] | 0.0604 | 0.2193 | 88.61 | 0.0725 | 0.2313 | 94.81 |
@@
Comparison with baselines: We first evaluate the performance of orientation-aligned object generation. As shown in the Table 1 and Table 2, our fine-tuned models, Wonder3D-OA and Trellis-OA, surpass most two-stage baselines in geometry and appearance on both GSO and Toys4k datasets. Note that our method is suboptimal in appearance on the GSO dataset compared to the VLM baseline. It is probably because the GSO dataset has more objects with irregular geometry and appearance, which is challenging for our fine-tuned generator to handle. We present more qualitative results in the Figure 5 and Figure 6. As shown in Figure 6, PCA cannot handle objects with different shape features and has difficulty in distinguishing the direction of estimated principal axes. VLM has trouble recognizing objects with unclear front-view features, like dolphins and boats. Orient Anything is still not that accurate, and typically rotates objects into leaning poses. In contrast, our method accurately aligns the orientation of the generated 3D models without obvious geometry and appearance degradation. Besides, our method can handle real world internet images and multi-objects images as shown in Figure 8 and Figure 8. Furthermore, due to the high geometry quality of our Objaverse-OA, Trellis-OA eliminates the production of plane-like geometries. Please refer to the supplementary materials Section B for more qualitative results.
@@ 6.3 Zero-Shot Object Orientation Estimation
| Toys4k [Stojanov2021UsingST] | Stick-like Obj. from ImageNet3D [Ma2024ImageNet3DTG] | |||
@@
| FSDetView [Xiao2020FewShotOD] (Few-shot) | 20.90 | 91.66 | 10.29 | 84.25 |
| Orient Anything [Wang2024OrientAL] (Vit-S) | 42.05 | 52.72 | 2.63 | 81.70 |
| Orient Anything [Wang2024OrientAL] (Vit-L) | 63.18 | 36.37 | 9.8 | 78.19 |
@@
Comparison with baselines: As shown in Table 3, without specific training, our method is already comparable with the SOTA orientation estimation method [Wang2024OrientAL] with the ViT-large architecture and surpasses [Wang2024OrientAL] with the ViT-small architecture by a large margin. Besides, our method can handle challenging stick-like objects collected from the real-world dataset while baselines all fail to work, which further demonstrates our utility in coping with long-tailed situations.
@@ 7 Conclusion and Limitation
8 Acknowledgement
This work is supported by NSFC under grant 62441223 and 62202418, and partially by Ant Group.
@@ A.1 Dataset Curation
Data Processing Pipeline. We first apply Vision Language Model (VLM) recognition across the entire dataset, followed by manual identification and correction of failure cases. Specifically, four images were rendered for each 3D object using an orthogonal camera setup. The front view was identified using the VLM (Gemini 2.0 [gemini]), and the objects were subsequently rotated to their canonical poses based on the recognition results. Note that objects recognized as without front-view orientation are automatically excluded from the dataset. After this initial step, the orientations of the processed objects were manually refined using Blender software.
@@ A.2 Experiment Implementations
Evaluation Data. We evaluate our method on three unseen datasets: GSO [Downs2022GoogleSO], Toys4k [Stojanov2021UsingST], and Imagenet3D [Ma2024ImageNet3DTG]. For the GSO dataset, we randomly collected 48 objects with clearly defined front views. Each object was rendered into four images, sampled from the upper hemisphere with azimuth angles uniformly distributed within [0∘, 360∘], polar angles within [0∘, 60∘], and rotation angles within [-30∘, 30∘]. For the Toys4k dataset, we randomly selected 439 objects across 47 categories. Each object was rendered into a single image using the same camera configuration as applied to the GSO dataset. For the Imagenet3D dataset, we focused on seven stick-like object categories, comprising a total of 204 objects. Please refer to Table 4 and Table 5 for details.
@@
Baselines. To implement the PCA-based baseline, we first calculate the three principal axes by eigen-decomposing the 3D models’ geometries. The object is then rotated to align these principal axes with the global coordinate system’s x-, y-, and z-axes. For the VLM-based baseline, we follow the same method used in the dataset curation, utilizing Gemini-2.0 [gemini] for orientation recognition. To implement the baseline based on Orient Anything [Wang2024OrientAL], we use the official checkpoint based on the ViT-large architecture and adopt its data augmentation module. The object orientation is estimated from an image rendered using a fixed camera, and the object is subsequently rotated according to the predicted orientation. For the pre-trained 3D generative models based on multi-view diffusion and 3D-VAE, we utilize the official Wonder3D++ and Trellis checkpoints. For the orientation estimation baseline based on FSDetView [Xiao2020FewShotOD], we restrict evaluation to the object categories supported by the method.
Metrics. For the calculation of Chamfer Distance, we don’t perform PCA for both predicted 3D models and GT models (except the PCA baseline), in order to measure the pose canonicalization performance. As for the calculation of CLIP and LPIPS, for the 3D-VAE backbone, we render four orthogonal views with the camera elevation angle of 0 for both the generated and GT 3D models, and computed the LPIPS and CLIP scores based on these renderings at matched camera poses, while for the Multi-view Diffusion backbone, we randomly sample views on a unit sphere for evaluation to avoid unfairness, since the orthogonal views are aligned with the camera setting of our Wonder3D-OA, but not aligned with the Wonder3D baselines.
@@ A.3 Method Implementations
Efficient Arrow-based Object Rotation Manipulation. For the arrow-based object rotation manipulation in the augmented reality application, we define this arrow-based interaction using a 2D start point and a 2D end point , forming a direction vector . Then we use Unidepth [Piccinelli2024UniDepthUM] to estimate the camera intrinsic and depth map. With camera intrinsic, we calculate the 3D camera rays and for the corresponding 2D points and . With the depth map, we calculate the plane via the least square method. After that, we calculate the 3D intersection points and between the corresponding camera rays , and the extracted plane . Finally, 3D direction vector can be calculated: . Since each generated model is aligned to a known forward vector , we can directly compute and apply the rotation that aligns with , allowing the object to be placed in the correct orientation without any manual adjustment. Note that the 3D start point is also the 3D location of the inserted 3D object, the 3D models need to rotate along to ensure verticality to the ground and the object size still needs to be set via VLM or user interaction, since our generated models share a normalized scale. Specifically, to automatically estimate the size of the object to insert, we can directly ask the VLM to estimate the common size of the object in the real world with the single-view image as input, since our method estimates metric depth with the scale aligned with the real world for insertion. Besides, to make the insertion results realistic, we set a plane as a shadow catcher and estimate the environment lighting using DiffusionLight [Phongthawee2023DiffusionLightLP]. For the arrow-based object rotation manipulation in the generic 3D software, we define this arrow-based interaction using a 3D start point and a 3D end point , forming a 3D direction vector . The rotation transformation is calculated by aligning with using Rodrigues’ formula and users can further rotate the objects along if needed.